Phantomdex:

P(A) = 0.35 and P(B.) = 0.42. If events A and B are independent, what is P(A∩B)? 0.83 0.77 0.15 0.07

Phantomdex:

(Ignore the dot in the P(B.) I just had to put there so there wouldn't be an emoji)

Phantomdex:

@toga

toga:

If events A and B are independent, then we can use the formula P(A∩B) = P(A) * P(B). Therefore, using the values given in the problem, we have: P(A∩B) = P(A) * P(B) = 0.35 * 0.42 = 0.147 The closest option is 0.15.

Phantomdex:

Thank you